# Next-Gen Interconnection Systems with Compute Express Link: a Comprehensive Survey

### v1

@@ Abstract
Interconnection is crucial for computing systems. However, the current interconnection performance between processors and devices, such as memory devices and accelerators, significantly lags behind their computing performance, severely limiting the overall performance. To address this challenge, Intel proposes Compute Express Link (CXL), an open industry-standard interconnection. With memory semantics, CXL offers low-latency, scalable, and coherent interconnection between processors and devices. This paper introduces recent advances in CXL-based interconnection systems with memory semantics. We classify the existing research into three categories: Pooling Memory, Distributed Shared Memory, and Unified Memory. Pooling Memory interconnects processors and memory, aims to address memory wall challenge. Distributed shared memory interconnects processors across nodes, aims to synchronize the cluster. Unified memory interconnects processors and accelerators, aims to enhance collaboration in heterogeneous computing systems. Finally, we discuss the future research and envision memory-centric computing with CXL.
@@
Compute Express Link, Interconnection, Memory Wall, Memory Expansion, Distributed Shared Memory, Unified Memory
@@ I Introduction
Interconnection is becoming a performance bottleneck in computing systems. A complex and heterogeneous computing system consists of Computing Units, such as CPU, GPU and TPU, Storage Units, Interconnections, such as PCIe and NVlink. The performance of computing units has grown 60,000 times in 20 years, yet the bandwidth of the interconnect has only grown 30 times during the same time[1]. Lagging interconnect performance results in the compute units needing to waste a lot of performance to reach synchronisation with the interconnect.
@@
In computing systems, interconnections can be categorized into three types: interconnections between processors and memory, interconnections between processors across nodes, and interconnections between heterogeneous accelerators. Interconnection bottleneck between the processors and memory result in Memory Wall, the pin-inefficiency of DDR memory limits the bandwidth of the processor to access the memory, as well as the memory capacity a processor can access[2]. Interconnection bottleneck between processors across nodes results in significant overhead for data synchronization in a distributed system. In the 1990s, distributed shared memory(DSM) established a memory abstraction of the global address space on distributed systems. However, the high latency of network communication made it difficult to maintain cache coherence across nodes. Interconnection bottleneck between heterogeneous accelerators leads to a lack of efficient collaboration. The root cause is the Non-Coherent Access from device to device and device to processor. Data has to be moved to device-attached memory before computation, resulting in a lot of data movement.
@@
To address these challenges, Intel proposed Compute Express Link (CXL)[3], [4], [5], an open industry-standard interconnection in 2019. CXL offers coherency and memory semantics with bandwidth that scales with PCIe bandwidth while achieving significantly lower latency than PCIe[2].
@@
CXL defines a set of protocols between processors and devices: CXL.io, CXL.mem, CXL.cache. The CXL.io protocol is based on PCIe. it is used for device discovery, status reporting, virtual-to-physical address translation, and direct memory access[2]. CXL.cache is used by devices to cache system memory. CXL.mem enables processors to access the device-attached memory as cachable memory, called Host-managed Device Memory(HDM), thus enabling a unified view(as shown in Fig 1) of the host between HDM and host memory.
Figure 1: The unified View of the host between HDM and host memory[6]
@@
Based on the combination of the protocols, CXL proposes three types of devices to address different scenarios. Type 1 devices are accelerators without memory (e.g., NICs), which implement only the CXL.io and CXL.cache protocols, using coherency semantics along with PCIe-style DMA transfers[2]. Type 2 devices are accelerators with memory (e.g., GPUs), which implement all CXL protocol and is able to cache each other’s memory with the processor. Type-3 devices implement only the CXL.io and CXL.mem protocols for memory bandwidth and capacity expansion. Compared to DDR memory, Type-3 devices are more cost, power and pin-efficient[2].
@@
This survey aims to fill the gaps left by previous survey, focusing on the latest CXL research and real-world use case, rather than delving deeply into the underlying hardware design. We research the latest work (to the end of 2024) and classify the existing research into three categories with memory semantics, based on the interconnection type.
The first part is Memory Expansion and Pooling, focusing on the interconnection between processors and memory devices(Type3 Device). Researchers not only discussed how to use a tiered CXL memory system[7, 8] on a single machine to optimize workload‘s performance[9, 10], but also attempted to build a distributed memory pooling system using CXL Switch proposed by CXL2.0[11],[12, 13]. They also try to use different storage techniques[14] to reduce costs, using near-data processing[15, 16] to avoid the overhead of data movement. The second part is Distributed Shared Memory, focusing on the interconnection between cross-host processors. CXL3.0 introduces cross-node memory sharing, which inspires researchers to use distributed shared CXL memory for cross-node communication[17, 18] and optimize the performance of in-memory workloads such as redis[19]. The last part is Unified Memory, focusing on the interconnection between processors and heterogeneous accelerator. CXL‘s hardware cache coherence allows heterogeneous devices to uniformly cache data on each other’s memory, which not only avoid unnecessary data copying[20, 21], but also provides more memory capacity for heterogeneous devices[22].
@@
The rest of the survey is organized as follows. Section II introduces the performance evaluation of existing CXL systems, the real world use cases, and the simulation platform for CXL hardware, providing performance guidance and platform support for future research. Section III introduces the tiered CXL memory system and memory pooling system based on CXL switch. Section IV introduces the researches on distributed shared CXL memory. Section V presents a discussion about how to use CXL to accelerate heterogeneous computing systems. In Section VI, we propose 5 main directions for future research. In Section VII, we introduce the memory-centric computing paradigm in the CXL era. Finally, Section VIII gives a conclusion. Figure 2 shows an overview of CXL based interconnection systems by interconnection type.
Figure 2: An Overview of CXL Based Interconnection Systems
@@ II Evaluation and Simulation
Before discussing the existing research, we first introduce the evaluation results of real CXL systems. Existing work [23, 24], [25, 26], [27] evaluates CXL from two perspectives: the basic performance and real-world use case. Based on the results, researchers not only can gain a deeper understanding of the performance characteristics of CXL hardware, but also can guide the construction of their own experimental environment. This section mainly introduces the results of [25, 26], [27]
@@
In 2023, researchers from the University of Michigan proposed TPP[7], a page placement approach for the tiered memory system which includes CXL memory. TPP provides the latency characteristics of CXL memory and other memory technologies, which have been widely referenced, as shown in the following table I.
TABLE I: Characteristics of Memory Technologies.
@@
In the same year, Pond[11] from Microsoft Azure not only measured the end-to-end latency of host direct access to CXL memory, but also stated the tradeoff between latency and scale. In Pond’s evaluation, in an 8-socket system, the latency of the host direct access to the CXL memory controller is 155ns, while in a 16-socket system, the latency of access to a more distant CXL memory controller through a retimer is 180ns. In larger systems (32-64 sockets), the latency of connecting through retimer + CXL switch will exceed 270ns.
The researchers from UIU and Intel[25] evaluated and compared the performance of real CXL hardware from three different vendors. Each device has different CXL IP (ASIC-based hard IP and FPGA-based soft IP) and DRAM technologies (DDR5-4800, DDR4-2400, and DDR4-3200). This work proposed several important conclusions:
@@
The researchers from ByteDance and Harvard[26] focused on the currently commercially available ASIC-based CXL Type3 devices, and conducted a more in-depth evaluation, proposing several new conclusions.
The micro-benchmark results show that the ASIC-based CXL Type3 device significantly outperforms the FPGA-based CXL Type3 device. Interestingly, the researchers found that the location of the CXL hardware is very important: When the CXL device is located on the local socket, its latency is around 256ns, which increases slightly with the ratio of read&write operations. When the CXL hardware is on a remote socket, the latency increases sharply to 485ns. Additionally, as the access bandwidth increases, the latencies of both CXL and DRAM gradually increase, and when reaching a certain threshold, the latency will increase significantly.
Jie Liu et al[27] also evaluates CXL hardware and make more in-depth conclusion. The result tells that the bandwidth of CXL hardware is much lower than regular memory, limited by PCIe (the test platform is PCIe Gen5), with CXL hardware able to sustain around 8 threads (regular memory is 28 threads), and the peak bandwidth of CXL hardware is only about 40% of regular memory.
@@
Jacob Wahlgren et al.[28]provide a CXL simulation platform and an analysis tool to identify the memory usage patterns of applications and their optimization opportunities. Their evaluation used two modes:
@@
    The first mode kept the total bandwidth consistent but changed the ratio of DRAM and CXL memory, using existing scientific computing applications[29] as the workload.
@@
    The second mode maintained a consistent memory ratio but varied the CXL memory bandwidth, using graphics applications as the workload.
The evaluation results showed that even with 75% of the data placed in the pooled memory, the performance impact on some applications was less than 10%. Furthermore, the study found that dynamically configuring high-bandwidth systems by increasing the number of CXL links can effectively support bandwidth-intensive applications, such as OpenFOAM based on unstructured grids.
The researchers from UIU and Intel[25] evaluates real cxl hardware on latency-sensitive applications(Redis[19] with YCSB[30], FIO[31], etc.) and throughput applications(DLRM inference[32], etc.). The results show a 10-82% performance degradation when running applications directly on CXL memory. Even with dynamic page migration techniques like TPP[7], it can still increase tail latency.
The researchers from ByteDance and Harvard[26] conducts detailed evaluations on four real-world workloads: KVS, Spark, ECS, and CPU-based LLM inference. The consistent conclusion is that the mixed use of CXL memory and DRAM(Memory Interleave) does cause some performance degradation, but it is better than using SSDs. Tiered memory strategies like TPP[7] can effectively improve application performance in the KVS scenario, but perform poorly in data analysis scenarios like Spark.
Jie Liu et al.[27] also evaluated CXL hardware on HPC applications. Similar to previous work, compute-intensive applications with low memory access can benefit the most from CXL, with relatively small performance degradation. Bandwidth-intensive applications like data analytics, however, will be greatly impacted.
@@
Jie Liu et al[27]. also discusses the impact of CXL on large language models (LLMs) in depth, from three perspectives. First, on tensor offloading, where objects are offloaded from GPU memory to CXL memory, but due to the lack of CXL-enabled accelerators, this has to go through a relatively long path (GPU-CPU-CXL), severely impacting performance. Second, based on ZeRO-Offload[33], it evaluates the impact of CXL on the LLM training, concluding that compared to regular memory, CXL causes performance degradation in training, as parameter computation requires frequent access to the slower CXL memory. Finally, based on FlexGen[34], it evaluates the impact of CXL on the LLM inference, finding that the Decode stage is well-suited for CXL as it revolves around the KV cache without intensive memory accesses, while the Prefill stage sees performance degradation due to both latency and bandwidth limitations of CXL.
@@ II-D The Simulation Platform
Since most of the CXL hardware are currently in the prototype stage, only a few products have been commercialized and put into practical application. These products are expensive and hard to get. Therefore, it is not easy for researchers to acquire CXL hardware. Fortunately, some simulation platform targeting CXL hardware has been proposed, with their help, researchers can simulate a system which equipped with CXL hardware. We will also introduce these simulation platform in this section.
@@
Jacob Wahlgren et al.[28] implemented a cxl memory simulation tool based on the libnuma library, using remote numa node to simulate CXL memory.
@@
Researchers from Samsung, Qirui Yang et al.[35], also based on the NUMA architecture, used remote NUMA node to simulate CXL hardware. Unlike other work, Yang attempted to build a hierarchical memory architecture to simulate not only DRAM memory but also NVMe SSD on the simulated CXL NUMA node. Applications or VMs prioritize accessing memory on CXL devices, and when CXL memory is exhausted, they will use the Linux kernel’s Swap mechanism to swap memory pages between CXL memory and SSD.
@@
Yiwei Yang from UCSC proposed CXLMemSim[36], a pure software-simulated tool for analyzing the performance characteristics of CXL memory. CXLMemSim not only can be used to simulate CXL memory, but also proposes an event-driven performance model that attaches to existing unmodified programs, dividing program execution into multiple phases and collecting performance monitoring events at the end of each phase. Based on these events, it calculates the execution time to evaluate the performance impact of CXL memory on the application.
@@
Researchers from IIT, Raja Gond et al., proposed emucxl[37], an emulation framework for CXL memory. emucxl provides a user-space library and a NUMA-based CXL simulation backend, allows developers and researchers to rapidly prototype, without the need to rebuild the CXL platform and access abstraction layers. The work also demonstrates the application of emucxl in two different use cases: direct application access and application middleware.
@@
Pouya et al. proposed the Memory Stress (Mess) framework[38], which provides a comprehensive evaluation of almost all memory systems, including the simulation and analysis of the CXL memory. In the paper, the authors used the Mess simulator to simulate the CXL memory based on the bandwidth-latency curves provided by the manufacturers. The authors have made the Mess framework’s source code publicly available, providing the community with a powerful tool for researching and utilizing these new memory technologies.
@@
Researchers from IIT, Amit Puri et al., proposed DRackSim[39], a simulation infrastructure for modeling large-scale disaggregated memory systems supporting CXL. DRackSim simulates multiple compute nodes, memory pools, local/global memory managers, and interconnects that support coherent memory access. Additionally, DRackSim integrates a modified version of DRAMSim2 to enable simulation of local and remote memory. DRackSim provides a powerful tool for researchers to evaluate and optimize new designs for large-scale disaggregated memory systems in data centers.
III Memory Expansion and Disaggregation
III-A CXL Based Memory Expansion

### v2

# Next-Gen Computing Systems with Compute Express Link: a Comprehensive Survey

@@ Abstract
Interconnection is crucial for computing systems. However, the current interconnection performance between processors and devices, such as memory devices and accelerators, significantly lags behind their computing performance, severely limiting the overall performance. To address this challenge, Intel proposes Compute Express Link (CXL), an open industry-standard interconnection. With memory semantics, CXL offers low-latency, scalable, and coherent interconnection between processors and devices. This paper introduces recent advances in CXL-based computing systems from single-machine to distributed. In single-machine systems, we classify existing research into two categories: Memory Expansion and Unified Memory. Memory Expansion focus on processors and memory, aims to address memory wall challenge. Unified memory focus on processors and accelerators, aims to enhance collaboration in heterogeneous computing systems. In distributed systems, we present how to build efficient disaggregation systems based on CXL infrastructure, enabling resource pooling and sharing. Finally, we discuss the future research and envision memory-centric computing with CXL.
@@
Compute Express Link, Interconnection, Memory Expansion, Disaggregation, Distributed Shared Memory, Unified Memory, Near-Memory Processing
@@ I Introduction
Interconnection is becoming a performance bottleneck in computing systems. A complex and heterogeneous computing system consists of Computing Units, such as CPU, GPU and TPU, Storage Units, Interconnections, such as PCIe and NVlink. The performance of computing units has grown 60,000 times in 20 years, yet the bandwidth of the interconnect has only grown 30 times during the same time[1]. Lagging interconnect performance results in the compute units needing to waste a lot of performance to reach synchronisation.
@@
In single-machine computing systems, interconnections can be categorized into two types: interconnections between processors and memory buffers, interconnections between processors and heterogeneous accelerators. The bottleneck between the processors and memory buffers result in Memory Wall, the pin-inefficiency of DDR memory limits the bandwidth of the processor to access the memory, as well as the memory capacity a processor can access[2]. The bottleneck between processors and heterogeneous accelerators leads to a lack of efficient collaboration. The root cause is the Non-Coherent Access from device to device and device to processor. Data has to be moved to device-attached memory before computation, resulting in a lot of data movement.
With the rapid growth in the scale of data center applications (e.g., Deep Learning, Big Data), single-machine computing systems are unable to meet application‘s demands in terms of computing power and storage capacity. Computing systems are gradually evolving towards distributed and large-scale. In distributed computing systems, Disaggregation[3] and Sharing[4] can effectively improve the efficiency and overall resource utilization. However, the current distributed interconnection is built on the basis of network (TCP, RDMA, etc.). Network‘s high latency greatly hinders resource disaggregation and sharing. The overhead of a computing unit accessing computing and storage resources on other nodes through the network is much higher than accessing local resources.
@@
To address these challenges, Intel proposed Compute Express Link (CXL)[5], [6], [7, 8], an open industry-standard interconnection in 2019. CXL offers coherency and memory semantics with bandwidth that scales with PCIe bandwidth while achieving significantly lower latency than PCIe[2].
@@
CXL defines a set of protocols between processors and devices: CXL.io, CXL.mem, CXL.cache. The CXL.io protocol is based on PCIe. it is used for device discovery, status reporting, virtual-to-physical address translation, and direct memory access[2]. CXL.cache is used by devices to cache system memory. CXL.mem enables processors to access the device-attached memory as cachable memory, called Host-managed Device Memory(HDM), thus enabling a unified view of the host between HDM and host memory.
Figure 1: Three types of CXL devices. The Type 1 device can directly cache host memory, the Type 2 device can directly cache each other’s memory with the host, and the host can directly cache memory on a Type 3 device.
@@
Based on the combination of three protocols, CXL proposes three types of devices to address different scenarios, as shown in figure 1. Type 1 devices are accelerators without memory (e.g., NICs), which implement only the CXL.io and CXL.cache protocols, using coherency semantics along with PCIe-style DMA transfers[2]. Type 2 devices are accelerators with memory (e.g., GPUs), which implement all CXL protocol and is able to cache each other’s memory with the processor. Type-3 devices implement only the CXL.io and CXL.mem protocols for memory bandwidth and capacity expansion. Compared to DDR memory, Type-3 devices are more cost, power and pin-efficient[2].
@@
Our survey aims to fill the gaps left by previous survey, focusing on the latest CXL research and real-world use case, rather than delving deeply into the underlying hardware design. We research the latest works (to the end of 2024) and present them with memory semantics, from single-machine to distributed system.
In single-machine computing system, we start with the interconnection between processors and memory buffers (Type3 Device), discussing how to use CXL memory for capacity expansion[9, 10] and optimize workload‘s performance[11, 12]. Then, in order to avoid the overhead of data movement between processors and memory buffers. we discuss the Near-Memory processing [13, 14], which focuses on how to offload computing functions to the memory side.
In addition to memory buffer, we also study the interconnection between processors and heterogeneous accelerators . CXL‘s hardware cache coherence allows devices to uniformly cache data on each other’s memory, which not only avoid unnecessary data copying [15, 16], but also provides more memory capacity for heterogeneous devices[17].
From single-machine to distributed system, CXL 2.0 proposes Memory Pooling with the CXL Switch. Our survey discusses how to decouple compute and memory resources from nodes and build a fully disaggregation system [18],[19, 20], enabling flexible and efficient resource utilization. We also focus on CXL based Distributed Shared Memory. In CXL 3.0, a memory module can be shared by multiple compute modules, which inspires researchers to use shared memory for distributed communication [21, 22] and optimizing the performance of in-memory workloads such as redis[23].
@@
The rest of the survey is organized as follows. Section II introduces the performance evaluation of existing CXL systems, the real world use cases, and the simulation platform for CXL hardware. Section III introduces the memory expansion with CXL Type3 Device. Section IV discusses how to avoid the costly data movement with near-memory computing. Section V presents a discussion about how to use CXL to accelerate heterogeneous computing systems. Section VI talks about CXL based disaggregated system. Section VII introduces the researches on distributed shared CXL memory. In Section VIII, we propose five main directions for future research. In Section IX, we introduce the memory-centric computing paradigm in the CXL era. Finally, Section X gives a conclusion. Table I shows an overview of CXL based computing systems from single-machine to distributed.
TABLE I: An Overview of CXL Based Computing Systems
Scale Classification Research Topics Reference
Single-machine and Distributed Evaluation and Simulation The Latency and Bandwidth. [9],[18],[24],[25],[26]
Use Case: Regular Applications. [27],[24],[25],[26]
Use Case: Large Language Models. [25],[26],[28],[29]
The Simulation Platforms [27],[30],[31],[32],[33],[34]
Single-machine Memory Expansion Tiered Memory System [9], [10], [35, 36], [37], [38]
Flash Memory as a Backend [39, 40],[41]
Application-Specific Optimizations [11, 12], [42], [14], [25, 43]
Near-Memory Processing Workload-Customized Near-Memory Process [44],[45],[46],[14]
Generalized Near-Memory Process [47], [48],[13],[49],[50]
Unified Memory Memory Expansion with CPU Relay [17],[51]
Unified Memory with Direct CXL Access [52],[53],[15],[16],[54],[55]
Distributed Memory Pooling The Architecture of Pooling System [19],[56],[20],[57],[58]
The Interconnection in Pooling System [59],[60],[61],[62],[63],[64]
The Cost of Pooling System [65],[18, 66]
Memory Sharing Distributed Shared Memory with CXL 3.0 [67],[22],[68]
Communication via Shared CXL Memory [22],[69],[21],[70]
Failure Tolerance [22],[71]
@@ II Evaluation and Simulation
Before discussing the existing research, we first introduce the evaluation results of real CXL systems. Existing work [72, 73], [24, 25], [26] evaluates CXL from two perspectives: the basic performance and real-world use case. Based on the results, researchers not only can gain a deeper understanding of the performance characteristics of CXL hardware, but also can guide the construction of their own experimental environment. This section mainly introduces the results of [24, 25], [26].
@@
In 2023, researchers from the University of Michigan propose TPP[9], a page placement approach for the tiered memory system which includes CXL memory. TPP provides the latency characteristics of CXL memory and other memory technologies, which have been widely referenced, as shown in the following table II.
TABLE II: Characteristics of Memory Technologies.
@@
In the same year, Pond[18] from Microsoft Azure not only measures the end-to-end latency of host direct access to CXL memory, but also states the tradeoff between latency and scale. In Pond’s evaluation, in an 8-socket system, the latency of the host direct access to the CXL memory controller is 155ns, while in a 16-socket system, the latency of access to a more distant CXL memory controller through a retimer is 180ns. In larger systems (32-64 sockets), the latency of connecting through retimer + CXL switch will exceed 270ns.
The researchers from UIU and Intel[24] evaluate and compared the performance of real CXL hardware from three different vendors. Each device has different CXL IP (ASIC-based hard IP and FPGA-based soft IP) and DRAM technologies (DDR5-4800, DDR4-2400, and DDR4-3200). This work proposes several important conclusions:
@@
The researchers from ByteDance and Harvard[25] focus on the currently commercially available ASIC-based CXL Type3 devices, and conduct a more in-depth evaluation, proposing several new conclusions.
The micro-benchmark results show that the ASIC-based CXL Type3 device significantly outperforms the FPGA-based CXL Type3 device. Interestingly, the researchers find that the location of the CXL hardware is very important: When the CXL device is located on the local socket, its latency is around 256ns, which increases slightly with the ratio of read&write operations. When the CXL hardware is on a remote socket, the latency increases sharply to 485ns. Additionally, as the access bandwidth increases, the latencies of both CXL and DRAM gradually increase, and when reaching a certain threshold, the latency will increase significantly.
Jie Liu et al[26] also evaluate CXL hardware and make more in-depth conclusion. The result tells that the bandwidth of CXL hardware is much lower than regular memory, limited by PCIe (the test platform is PCIe Gen5), with CXL hardware able to sustain around 8 threads (regular memory is 28 threads), and the peak bandwidth of CXL hardware is only about 40% of regular memory.
@@
Jacob Wahlgren et al.[27]provide a CXL simulation platform and an analysis tool to identify the memory usage patterns of applications and their optimization opportunities. Their evaluation uses two modes:
@@
    The first mode keeps the total bandwidth consistent but changes the ratio of DRAM and CXL memory, using scientific computing applications[74] as the workload.
@@
    The second mode maintains a consistent memory ratio but varies the CXL memory bandwidth, using graphics applications as the workload.
The evaluation results show that even with 75% of the data placed in the pooled memory, the performance impact on some applications is less than 10%. Furthermore, the study finds that dynamically configuring high-bandwidth systems by increasing the number of CXL links can effectively support bandwidth-intensive applications, such as OpenFOAM based on unstructured grids.
Researchers from UIU and Intel[24] evaluate real cxl hardware on latency-sensitive applications(Redis[23] with YCSB[75], FIO[76], etc.) and throughput applications(DLRM inference[77], etc.). The results show a 10-82% performance degradation when running applications directly on CXL memory. Even with dynamic page migration techniques like TPP[9], it can still increase tail latency.
Researchers from ByteDance and Harvard[25] conduct detailed evaluations on four real-world workloads: KVS, Spark, ECS, and CPU-based LLM inference. The consistent conclusion is that the mixed use of CXL memory and DRAM(Memory Interleave) does cause some performance degradation, but it is better than using SSDs. Tiered memory strategies like TPP[9] can effectively improve application performance in the KVS scenario, but perform poorly in data analysis scenarios like Spark.
Jie Liu et al.[26] also evaluate CXL hardware on HPC applications. Similar to previous work, compute-intensive applications with low memory access can benefit the most from CXL, with relatively small performance degradation. Bandwidth-intensive applications like data analytics, however, will be greatly impacted.
@@
Jie Liu et al[26]. also discuss the impact of CXL on large language models (LLMs) in depth, from three perspectives. First, tensor offloading, where objects are offloaded from GPU memory to CXL memory, but due to the lack of CXL-enabled accelerators, this has to go through a relatively long path (GPU-CPU-CXL), severely impacting performance. Second, based on ZeRO-Offload[28], they evaluate the impact of CXL on the LLM training, concluding that compared to regular memory, CXL causes performance degradation in training, as parameter computation requires frequent access to the slower CXL memory. Finally, based on FlexGen[29], they evaluate the impact of CXL on the LLM inference, finding that the Decode stage is well-suited for CXL as it revolves around the KV cache without intensive memory accesses, while the Prefill stage sees performance degradation due to both latency and bandwidth limitations of CXL.
@@ II-D The Simulation Platform
Since most of the CXL hardware are currently in the prototype stage, only a few products have been commercialized. Therefore, it is not easy for researchers to acquire CXL hardware. Fortunately, some simulation platforms targeting CXL hardware has been proposed, with their help, researchers can simulate a system which equipped with CXL hardware. In this section, we will introduce these simulation platforms.
@@
Jacob Wahlgren et al.[27] implement a cxl memory simulation tool based on the libnuma library, using remote numa node to simulate CXL memory.
@@
Researchers from Samsung, Qirui Yang et al.[30], also based on the NUMA architecture, use remote NUMA node to simulate CXL hardware. Unlike other work, Yang attempts to build a tiered memory architecture to simulate not only DRAM memory but also NVMe SSD on the simulated CXL NUMA node. Applications or VMs prioritize accessing memory on CXL devices, and when CXL memory is exhausted, they will use the Linux kernel’s Swap mechanism to swap memory pages between CXL memory and SSD.
@@
Yiwei Yang from UCSC proposes CXLMemSim[31], a pure software-simulated tool for analyzing the performance characteristics of CXL memory. CXLMemSim not only can be used to simulate CXL memory, but also proposes an event-driven performance model that attaches to existing unmodified programs, dividing program execution into multiple phases and collecting performance monitoring events at the end of each phase. Based on these events, it calculates the execution time to evaluate the performance impact of CXL memory on the application.
@@
Researchers from IIT, Raja Gond et al., propose emucxl[32], an emulation framework for CXL memory. emucxl provides a user-space library and a NUMA-based CXL simulation backend, allows developers and researchers to rapidly prototype, without the need to rebuild the CXL platform and access abstraction layers. The work also demonstrates the application of emucxl in two different use cases: direct application access and application middleware.
@@
Pouya et al. propose the Memory Stress (Mess) framework[33], which provides a comprehensive evaluation of almost all memory systems, including the simulation and analysis of the CXL memory. They simulate the CXL memory based on the bandwidth-latency curves provided by the manufacturers. They have made the Mess framework’s source code publicly available, providing the community with a powerful tool for researching and utilizing these new memory technologies.
@@
Researchers from IIT, Amit Puri et al., propose DRackSim[34], a simulation infrastructure for modeling large-scale disaggregated memory systems supporting CXL. DRackSim simulates multiple compute nodes, memory pools, local/global memory managers, and interconnects that support coherent memory access. Additionally, DRackSim integrates a modified version of DRAMSim2 to enable simulation of local and remote memory. DRackSim provides a powerful tool for researchers to evaluate and optimize new designs for large-scale disaggregated memory systems in data centers.
III Memory Expansion